OpenStudy (anonymous):

Can someone please explain to me how to find an equation of the line containing the given points (1/6,-1/3) and (5/6,3)

OpenStudy (amistre64):

yes :)

OpenStudy (amistre64):

1) find out the "slope" between the points by subtracting one from the other

OpenStudy (amistre64):

x,y (5/6 , 3 ) - (1/6, -1/3) ------------- (4/6,8/3) slope = y/x = (8/3)/(4/6) and that can be simplified

OpenStudy (anonymous):

:D y-y1=m(x1-x)

OpenStudy (amistre64):

8/3 8 6 --- = -- * -- = 2*2 = 4 4/6 4 3 slope is = 4

OpenStudy (amistre64):

the equation of a line is y = "slope" * x + b; and it needs to contain at least one of the points given :) perferably both; but we have the slope, so we only need one of them to work it out

OpenStudy (amistre64):

y = mx + b; (x=5/6, y=3) 3 = (4)(5/6) + b ; solve for b 3 = 20/6 +b 3 - 10/3 = b -1/3 = b the equation of the line is: y = 2x - 1/3 which can be written also as... 2x -y = 1/3

OpenStudy (amistre64):

.... arg!! typoed it lol

OpenStudy (amistre64):

y = 4x - (1/3) -4x +y = -1/3 4x -y = 1/3 are all equal equations for this line